OpenStudy (anonymous):

What is the simplified form of 4x^2-20x+25/2x-5

OpenStudy (anonymous):

and what are the restrictions plz explain so i can do the rest of examples

OpenStudy (anonymous):

that doesnt help i wasnt asking for that i want an explanation

pooja195 (pooja195):

oh

OpenStudy (anonymous):

yea so can you explain

pooja195 (pooja195):

so u want to kno how to simplify that right

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

like evertyhing how to get the restritiond

OpenStudy (anonymous):

restrictions*

pooja195 (pooja195):

yeah im not sure @ZeHanz any ideas

OpenStudy (anonymous):

how did you get the answer you gave me

OpenStudy (anonymous):

im that type of person where i dont want food sped to me on a silver platter

OpenStudy (anonymous):

i want a explanation

OpenStudy (anonymous):

plz?

OpenStudy (zehanz):

If you want to simplify a fraction, there has to be some common factor in numerator and denominator. The denominator has only 2x-5 so that is the only factor. It is a good idea to factor the numerator then. Hopefully, 4x^2-20x+25 will contain 2x-5 as a factor, so we can simplify!

OpenStudy (anonymous):

ok the restrictions

OpenStudy (zehanz):

What did you have in mind?

OpenStudy (anonymous):

know i wanna know how to get that i dont understand that

OpenStudy (anonymous):

no*

OpenStudy (zehanz):

Or, to be more specific, in the kind of form you are dealing with, what kind of restriction could you think of?

OpenStudy (anonymous):

x not equal to 5/2

OpenStudy (zehanz):

There you go!

OpenStudy (anonymous):

i mean negative

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

i meant x not equal to -5/2

OpenStudy (zehanz):

Stay positive on this one, dear!

OpenStudy (anonymous):

just explain

OpenStudy (anonymous):

im on negative

OpenStudy (zehanz):

2*(-5/2)-5=-5-5=-10 >> dividing by -10 is not a problem.

OpenStudy (anonymous):

i just dont understand the restrictions part

OpenStudy (anonymous):

look here are the answers choices

OpenStudy (zehanz):

We don't need choices, we need to think. Why would there be any restrictions? Let's look at the numerator: can you calculate 4x²-20x+25, no matter what x is?

OpenStudy (anonymous):

theres restrictions on my problem and i dont even know what the heck that is

OpenStudy (anonymous):

thats the part thats confusing me can you kindly explain that part

OpenStudy (zehanz):

I'm trying: Why would there be any restrictions? Let's look at the numerator: can you calculate 4x²-20x+25, no matter what x is?

OpenStudy (anonymous):

yes its 2x-5

OpenStudy (zehanz):

No, the numerator is 4x²-20x+25. Can you calculate 4x²-20x+25, no matter what x is?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

that doesnt make sense isnt there suppose to be =0

OpenStudy (anonymous):

4x^2-20x+25=0 can you calculate that you meant

OpenStudy (zehanz):

Why? All you need to do is the following: 1. calculate 4x². can do that for every x! 2. subtract 20x. no problem 3. add 20. no problem. See? You can put ANY number in 4x²-20x+25, no matter what x is! There will always be an outcome. Denominator 2x-5: exactly the same: we can double the value of ANY number and then subtract 5. Yet, there IS a value of x may may NOT use if we have to do with 4x²-20x+25 and 2x-5 what has to be done: we have to divide 4x²-20x+25 by 2x-5. When can you not divide?

OpenStudy (anonymous):

ok you get 2x-5

OpenStudy (anonymous):

when you divide but still your not explaining me how to get restrictions

OpenStudy (zehanz):

I am doing it all the time, but you have to read carefully what I wrote. We choose some (any) x. All we have now is two numbers: the value of 4x²-20x+25 and the value of 2x-5. Now we need to divide the two. You must ask yourself this question: can I always divide two numbers, no matter what the numbers are?

OpenStudy (anonymous):

yes

OpenStudy (zehanz):

So tell me how much is 3/0

OpenStudy (anonymous):

0

OpenStudy (zehanz):

If that was true, then I could check this calculation in this way: 18/3=6, so 6*3 = 18, right? So if 3/0 = 0, then 0*0=3?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

0*0 doesnt equal 3

OpenStudy (zehanz):

Of course not! It is worse than that: if 3/0=a, then a*0=3. There isn't a number to be found that multiplied with 0 would give 3! So: 3/0 cannot be done. That's the restriction.

OpenStudy (zehanz):

You always have to remember this very important fact: You cannot divide by zero!

OpenStudy (anonymous):

ok can you use that example in the problem i gave u and maybe i will understand it

OpenStudy (zehanz):

In the problem you have to divide by 2x-5. If you set x such that 2x-5 becomes 0, you have a problem, because you cannot divide by 0. So when IS 2x-5=0? Well, then 2x=5, and x=5/2. So the one and only number that you may NOT use in this division, is x =5/2, because it would lead to division by 0.

OpenStudy (anonymous):

and you can divide by zero 5/0=0

OpenStudy (anonymous):

0/5=0

OpenStudy (zehanz):

dividing 0 by 5 is no problem. Check: 0/5=0, because 5*0=0 OK

OpenStudy (anonymous):

whats 0/5

OpenStudy (anonymous):

its 0 isnt it

OpenStudy (zehanz):

Yup!

OpenStudy (anonymous):

ok your still not making sense

OpenStudy (anonymous):

did you explain it how the problem i gave u

OpenStudy (zehanz):

Well, I am: 0/5=0 because 5*0=0. Seems senible enough!

OpenStudy (anonymous):

ok i understand that but how does that help me with my problem

OpenStudy (zehanz):

A lot. In your problem you have to do 2 things: 1. Simplify the fraction. (I gave you a hint way up in this topic) 2. State restrictions on the variable. We found this restriction: x may not be 5/2. So you still have to do 1., but your questions focused on 2. all the time...

OpenStudy (anonymous):

im asking how to get number 2

OpenStudy (anonymous):

and your giviving me examples off other things

OpenStudy (zehanz):

We have done that. Restriction found: x may not be 5/2.

OpenStudy (anonymous):

0/5=0 3/0=0 etc.

OpenStudy (anonymous):

how

OpenStudy (anonymous):

your just giving me the answer explain

OpenStudy (anonymous):

plz?

OpenStudy (anonymous):

the restrictions

OpenStudy (anonymous):

i dont get

OpenStudy (anonymous):

i get everything else

OpenStudy (anonymous):

can you put that in math form plx

OpenStudy (anonymous):

oh

OpenStudy (anonymous):

and that would be x not equal to 5/2

OpenStudy (anonymous):

I live in USA, but im jamaican

OpenStudy (anonymous):

can you explain

OpenStudy (anonymous):

read the problem cause im getting different answers

OpenStudy (anonymous):

which is

OpenStudy (anonymous):

x not equal to -5/2

OpenStudy (anonymous):

can i just solve for x and set =0

OpenStudy (anonymous):

its -2x-5=0

OpenStudy (anonymous):

but if i solve i still get 5/2

OpenStudy (anonymous):

x not eqqual to 5./2

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

after an hour i finally get it

OpenStudy (anonymous):

Thank you so Muhc

OpenStudy (anonymous):

ok